# Sharp Bounds for Multiple Models in Matrix Completion

### v2

@@ Abstract
In this paper, we demonstrate how a class of advanced matrix concentration inequalities, introduced in Brailovskaya and van Handel [2024], can be used to eliminate the dimensional factor in the convergence rate of matrix completion. This dimensional factor represents a significant gap between the upper bound and the minimax lower bound, especially in high dimension. Through a more precise spectral norm analysis, we remove the dimensional factors for five different estimators in various settings, thereby establishing their minimax rate optimality.
@@ 1 Introduction
In particular, many studies focus on the convergence rate of the nuclear norm penalization methods [Koltchinskii et al., 2011, Negahban and Wainwright, 2012, Klopp, 2014, Yu et al., 2024]. However, the minimax lower bound, obtained in Koltchinskii et al. [2011] and Negahban and Wainwright [2012], does not include a logarithmic dimension factor that always appears in the upper bound of the convergence rate. This discrepancy between the two bounds becomes more pronounced in high-dimensional settings, highlighting a potential gap in the existing theoretical understanding. Due to this issue, previous works have had to qualify their results by stating “our estimator is minimax optimal, up to a logarithmic factor”. In this paper, we use a class of sharp matrix concentration inequalities to remove the logarithmic dimension factor. After our work, the qualification will no longer be necessary.
@@
Let $A_{0}\in\mathbb{R}^{m_{1}\times m_{2}}$ be the unknown matrix we aim to recover. We observe independent samples $\{(X_{i},Y_{i})\}_{i=1}^{\nu}$ from the model
$Y_{i}=\langle X_{i},A_{0}\rangle+\xi_{i},\quad i=1,\ldots,\nu,$ (1)
where $X_{i}\in\mathbb{R}^{m_{1}\times m_{2}}$, $\langle X_{i},A_{0}\rangle=\text{tr}(X_{i}^{T}A_{0})$, and $\text{tr}(B)$ denotes the trace of matrix $B$. The noise variables $\xi_{i}$ are independent and identically distributed. The sampling matrices $X_{i}\in\mathbb{R}^{m_{1}\times m_{2}}$ take values from $\{e_{l}(m_{1})e_{k}^{T}(m_{2}),1\leq l\leq m_{1},1\leq k\leq m_{2}\}$, where $e_{j}(m_{1}),e_{k}(m_{2})$ denote standard basis vectors in $\mathbb{R}^{m_{1}},\mathbb{R}^{m_{2}}$, respectively. This implies that each $X_{i}$ only has one non-zero entry, indicating the location of the sampled entry.
@@
$P_{jk}=P(X_{i}=e_{j}(m_{1})e_{k}^{T}(m_{2})).$
Additionally, $X_{i}$ and $\xi_{i}$ are assumed to be independent. In this paper, we consider that $\text{rank}(A_{0})\leq r$ or $A_{0}$ is the sum of a low rank matrix and a sparse matrix, and $\|A_{0}\|_{\infty}\leq a$ where $\|A\|_{\infty}=\max_{1\leq i\leq m_{1},1\leq j\leq m_{2}}|a_{ij}|$ is the entrywise maximum norm of $A=(a_{ij})$. The total sample size $\nu=n{\rm\leavevmode\nobreak\ or\leavevmode\nobreak\ }N$ depending on the settings.
@@
The matrix completion problem is often addressed by a penalized regression. Assuming $A_{0}$ is of low rank, a natural choice for the penalty is the rank of a matrix. However, the rank penalty is non-convex and makes the corresponding program difficult to solve. The nuclear norm $\|\cdot\|_{1}$, namely the sum of all the singular values of a matrix, is usually used as a convex relaxation for the rank. There are many classical works on nuclear norm penalization methods under different assumptions. Assuming the noise has exponential decay, Koltchinskii et al. [2011] consider the case that the sampling distribution is known; Negahban and Wainwright [2012] study the situation where the sampling distribution is unknown and the row indices and the column indices of the observations are sampled independently; Klopp [2014] explores the problem for general sampling distributions. Consider the classical nuclear norm penalized least squares estimator
$\hat{A}=\operatorname*{arg\,min}_{\|A\|_{\infty}\leq a}\left\{\dfrac{1}{n}\sum_{i=1}^{n}(Y_{i}-\langle X_{i},A\rangle)^{2}+\lambda\|A\|_{1}\right\}.$
@@
$\dfrac{\|\hat{A}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\apprle\log(m_{1}+m_{2})\dfrac{r\max(m_{1},m_{2})}{n},$
@@
$\inf_{\hat{A}}\sup_{\mathcal{A}}P\left(\dfrac{\|\hat{A}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\apprge\dfrac{r\max(m_{1},m_{2})}{n}\right)\geq C,$
@@
Recently, Bandeira et al. [2023] and Brailovskaya and van Handel [2024] introduce new matrix concentration inequalities, which establish sharper bounds for certain important random matrices. We showcase how the new bounds can be leveraged in the matrix completion problem to drop the logarithmic dimension factor. Specifically, We consider five estimators across three different settings as follows:
  • First setting: the unknown matrix is of low rank and the noise is sub-Gaussian. This is the most classical setting. Two estimators are studied in Klopp [2014], one for the case of known variance and another for the case of unknown variance.
  • Second setting: the unknown matrix is of low rank and the noise is heavy tailed. The estimator in Yu et al. [2024] is designed for the case that the noise is only assumed to have finite second moment.
  • Third setting: the unknown matrix is corrupted and the noise is sub-Gaussian. Two estimators are studied under different corruption models in Klopp et al. [2017]: entrywise corruption and columnwise corruption.
All the previous results on the convergence rates of the above five estimators involve the logarithmic dimension factor. We remove the dimensional factor for all the five estimators, thereby proving their minimax rate optimality.
The remainder of the paper is organized as follows. Continuing in this section, we introduce the notation. In Section 2, we present our main results. We begin by discussing a classical matrix completion model and demonstrating how to eliminate the dimension factor. We then introduce additional examples and results. We give some concluding remarks in Section 3. Section 4 is dedicated to the proofs.
@@
To present our results, we summarize the notations used throughout this paper.
@@
  • Let $A=(a_{ij})$, for $1\leq i\leq m_{1},1\leq j\leq m_{2}$. Let $\{\sigma_{l}\}_{l=1}^{m}$ denote ordered singular values of $A$: $\sigma_{1}\geq\sigma_{2}\geq\cdots\geq\sigma_{m}$. $\|A\|_{F}=\sqrt{\sum_{i=1}^{m_{1}}\sum_{j=1}^{m_{2}}a_{ij}^{2}}$, $\|A\|_{1}=\sum_{l=1}^{m}\sigma_{l}$, $\|A\|=\sigma_{1}$, $\|A\|_{\infty}=\max_{i\leq m_{1},j\leq m_{2}}|a_{ij}|$, $\|A\|_{1,*}=\sum_{i,j}|a_{ij}|$. With $A^{k}$ denoting the $k^{th}$ column of A, define $\|A\|_{2,1}=\sum_{k=1}^{m_{2}}\|A^{k}\|_{2}$, where $\|\cdot\|_{2}$ is the usual $l_{2}$-norm. $\|A\|_{w(F)}=\sqrt{\sum_{i=1}^{m_{1}}\sum_{i=1}^{m_{2}}a^{2}_{ij}P_{ij}}$ is a weighted Frobenius norm associated with the distribution of the sampling matrices $X_{i}$.
@@
    $P(\epsilon_{i}=1)=P(\epsilon_{i}=-1)=\frac{1}{2}.$
@@ 2 Main Results
In this section, we present new upper bounds for the five estimators mentioned in Section 1. In particular, compared to the existing bounds, the new bounds do not involve the dimension factor $\log d$ and hence match with the minimax lower bounds (under mild conditions).
2.1 A warming up example
We start with the most classical setting, where the noise in model (1) has exponentially decaying tails [Koltchinskii et al., 2011, Negahban and Wainwright, 2012, Klopp, 2014].
@@ 2.1 A warming up example
The noise variables $\xi_{i}$ satisfy $E[\xi_{i}]=0$, $E[\xi^{2}_{i}]=\sigma^{2}$ and
$E\exp(t\xi_{i}/\sigma)\leq\exp(t^{2}/L_{1}),\forall t\in\mathbb{R},$
where $L_{1}\geq 0$ is a constant.
@@
Under the assumption of sub-Gaussian tailed noise, a classical estimator for $A_{0}$ is:
$\hat{A}=\operatorname*{arg\,min}_{\|A\|_{\infty}\leq a}\left\{\dfrac{1}{n}\sum_{i=1}^{n}(Y_{i}-\langle X_{i},A\rangle)^{2}+\lambda\|A\|_{1}\right\}.$ (2)
We now present our new result:
Theorem 1.
Let Assumptions 1 and 2 hold. Assume that $n\geq C_{1}m\log^{6}d\cdot\log^{2}n$. There exists a constant $C_{2}$ such that if we take
$\lambda=C_{2}\sigma\sqrt{\frac{1}{nm}},$
with probability at least $1-\frac{2}{d}$, the following bound holds
$\dfrac{\|\hat{A}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\leq\dfrac{C_{3}\mu^{2}\max(\sigma^{2},a^{2})rM}{n}.$ (3)
Here, $C_{1}$ is a universal constant, and $C_{2},C_{3}$ depend on the constants $L_{1},L_{2},L_{3}$ from Assumptions 1-2.
To compare with existing bounds, we mention one representative result from Klopp [2014]:
Theorem 2 (Theorem 7 in Klopp [2014]).
Let Assumption 1 and Conditions (1)-(3) in Assumption 2 hold. Assume that $n\geq C_{1}m\log^{2}d$. If we take
$\lambda=C_{2}\sigma\sqrt{\dfrac{\log d}{nm}},$
for a suitable constant $C_{2}$, the following bound holds with probability $1-\frac{3}{d}$:
$\dfrac{\|\hat{A}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\leq C_{3}\max\left\{\dfrac{\mu^{2}\max(a^{2},\sigma^{2})rM\log d}{n},\mu a^{2}\sqrt{\frac{\log d}{n}}\right\}.$
Here, $C_{1}$ is a universal constant, and $C_{2},C_{3}$ depend on the constants $L_{1},L_{2}$ from Assumptions 1-2.
Comparing Theorems 1 and 2, we make several remarks:
  1. 1.
    Our new result eliminates the $\log d$ factor in the convergence rate. Koltchinskii et al. [2011] provides the following minimax lower bound in the case of gaussian noise and uniform sampling distribution:
    $\inf_{\hat{A}}\sup_{A_{0}}P\left(\dfrac{\|\hat{A}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\geq C_{1}\dfrac{\min(\sigma^{2},a^{2})rM}{n}\right)\geq C_{2},$ (4)
  2. 2.
    Our analysis reveals a more precise choice for the tuning parameter $\lambda$: the optimal choice for $\lambda$ should be of the order $O(\sqrt{1/(nm)})$ instead of $O(\sqrt{(\log d)/(nm)})$.
  3. 3.
    The new bound requires a slightly stronger condition on $n$, as $n\geq Cm\log^{6}d\log^{2}n$ is needed rather than $n\geq Cm\log^{2}d$. The $\log n$ term arises from the truncation of the sub-Gaussian noise, in order to apply the matrix concentration inequalities from Brailovskaya and van Handel [2024].
  4. 4.
    Compared to the sampling distribution assumption in Klopp [2014], Condition (4) of Assumption 2 is an additional condition we impose. But as we explained after Assumption 2, this is a very mild condition.
  5. 5.
    The convergence rate in Theorem 2 takes a form of the maximum between two terms. In the high-dimensional regime where $n\leq m_{1}m_{2}$, it is straightforward to verify that the first term–the desirable term–is the larger one. However, in less challenging scenarios in which $n$ is sufficiently large, the second term dominates. This limitation arises from a Frobenius norm based peeling argument that Klopp [2014] used to prove restricted strong convexity. In contrast, we have adopted a different peeling scheme introduced in Wainwright [2019], which peels certain matrix space by the infinity norm and the nuclear norm. As a result, our new bound does not include the nuisance term $O(\sqrt{(\log d)/n})$.
2.1.1 How to remove the $\log d$ factor
In this section, we discuss where the dimensional factor $\log d$ arises and how to eliminate it. Denote $L(A)=\frac{1}{n}\sum_{i=1}^{n}(Y_{i}-\langle X_{i},A\rangle)^{2}$. The high-dimensional analysis of (2) includes two basic steps [Negahban et al., 2012]. The first step involves selecting $\lambda$ such that
$\lambda\geq C\|\nabla L(A_{0})\|=2C\Big{\|}\dfrac{1}{n}\sum_{i=1}^{n}\xi_{i}X_{i}\Big{\|},$ (5)
to ensure that $\hat{A}-A_{0}$ lies in a nearly low rank space $\Omega$. In the second step, we establish strong convexity of $L(A)$ over $\{A:A-A_{0}\in\Omega\}$, namely,
$L(\hat{A})-L(A_{0})-\langle\nabla L(A_{0}),\hat{A}-A_{0}\rangle=\frac{1}{n}\sum_{i=1}^{n}\langle X_{i},\hat{A}-A_{o}\rangle^{2}\geq C\|\hat{A}-A_{0}\|^{2}_{w(F)}-\mathcal{E},$
where $\mathcal{E}$ represents a tolerance term and $\|A\|^{2}_{w(F)}=E\langle X_{i},A\rangle^{2}$. To achieve this, we shall obtain bounds on the following empirical processes
$Z=\sup_{A-A_{0}\in\Omega}\left|\dfrac{1}{n}\sum_{i=1}^{n}\langle X_{i},A-A_{0}\rangle^{2}-\|A-A_{0}\|^{2}_{w(F)}\right|.$
Bounding this type of empirical processes, after applying Talagrand concentration and symmetrization & contraction, invokes a bound on
$E\left\|\dfrac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}\right\|.$ (6)
With the two steps completed, we can use some standard non-asymptotic arguments to obtain upper bounds for the convergence rate in averaged Frobenius norm. Specifically, Lemma 14 reveals how (5) and (6) influence the error rate: with probability at least $1-\frac{1}{d}$, it holds
$\dfrac{\|\hat{A}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\apprle\mu^{2}\left(\lambda^{2}+a^{2}E^{2}\left\|\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}\right\|\right)rm_{1}m_{2}.$ (7)
Traditional matrix concentration inequalities (e.g. Theorem 6.17 and Example 6.18 in Wainwright [2019]; Lemmas 5 and 6 in Klopp [2014]), yield
$\left\|\frac{1}{n}\sum_{i=1}^{n}\xi_{i}X_{i}\right\|\leq C_{1}\sigma\sqrt{\frac{\log d}{nm}}$
holds with high probability and
$E\left\|\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}\right\|\leq C_{2}\sqrt{\frac{\log d}{nm}}.$
Plugging $\lambda=C_{1}\sigma\sqrt{\frac{\log d}{nm}}$ and $E\left\|\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}\right\|\leq C_{2}\sqrt{\frac{\log d}{nm}}$ into (7) leads to the dimension factor dependent result in Theorem 2.
To remove the $\log d$ factor, we refine the evaluation of the two key quantities $\|\frac{1}{n}\sum_{i=1}^{n}\xi_{i}X_{i}\|$ and $E\|\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}\|$. Recently, Bandeira et al. [2023] introduce new matrix concentration inequalities for random matrices with Gaussian entries using novel mathematical tools. Later, Brailovskaya and van Handel [2024] extend the results for more general distributions. This type of concentration inequalities fully take the non-commutativity of random matrices into account and provide sharper bounds. For instance, their results on the expectation of the spectral norm of a random matrix $X\in\mathbb{R}^{m_{1}\times m_{2}}$ typically take the form:
$E\|X\|\apprle\max(\|EX^{T}X\|,\|EXX^{T}\|)^{1/2}+v(X)\log^{\beta}d,$
where $\max(\|EX^{T}X\|,\|EXX^{T}\|)^{1/2}$ is the desirable part and $\beta$ is a positive constant. The term $v(X)$ is a statistic that captures the non-commutativity of the random matrix $X$ and is often smaller than $\max(\|EX^{T}X\|,\|EXX^{T}\|)^{1/2}$ in many situations. If we are given that
$v(X)\log^{\beta}d\leq\max(\|EX^{T}X\|,\|EXX^{T}\|)^{1/2},$
the bound is free of the dimensional factor $\log^{\beta}d$. Our computations confirm that this holds for $\|\frac{1}{n}\sum_{i=1}^{n}\xi_{i}X_{i}\|$ and $E\|\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}\|$. Assuming some mild conditions, we can obtain
$P\left(\left\|\frac{1}{n}\sum_{i=1}^{n}\xi_{i}X_{i}\right\|\geq C_{1}\sigma\sqrt{\dfrac{1}{nm}}\right)\leq\dfrac{C_{2}}{d},$
and
$E\left\|\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}X_{i}\right\|\leq C\sqrt{\dfrac{1}{nm}}.$
As a result, the $\log d$ factor is successfully removed. See Lemmas 12 and 13 for more details.
A similar issue appears in the existing results for the other four estimators. The advanced inequalities in Brailovskaya and van Handel [2024] can also be adapted to resolve the issue. We present sharper bounds for the remaining estimators in Sections 2.2-2.4.
2.2 Matrix completion with unknown variance of the noise
$\hat{A}_{S}=\operatorname*{arg\,min}_{\|A\|_{\infty}\leq a}\left\{\sqrt{\dfrac{1}{n}\sum_{i=1}^{n}(Y_{i}-\langle X_{i},A\rangle)^{2}}+\lambda\|A\|_{1}\right\}.$ (8)
We obtain a new result for (8).
Theorem 3.
Let Assumptions 1 and 2 hold. Let $n\geq C_{1}m\log^{6}d\log^{2}n$. If we take
$\lambda=C_{2}\sqrt{\frac{1}{nm}},$
with probability at least $1-\frac{2}{d}-2\exp(-C_{3}n)$, the following bound holds
$\dfrac{\|\hat{A}_{S}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\leq\dfrac{C_{4}\mu^{2}\max(\sigma^{2},a^{2})rM}{n},$ (9)
provided that $n\geq\frac{C_{2}}{10}\mu rM$. Here, $C_{1}$ is a universal constant, $C_{3}$ depends on $L_{1}$, and $C_{2},C_{4}$ depend on $L_{1},L_{2},L_{3}$, where $L_{i}$’s are constants from Assumptions 1-2.
We can compare the above bound with the one obtained in Klopp [2014].
Theorem 4 (Theorem 10 in Klopp [2014]).
Let Assumption 1 and Conditions (1)-(3) in Assumption 2 hold. If we take
$\lambda=C_{1}\sqrt{\frac{\log d}{nm}},$
the following upper bound holds with probability $1-\frac{3}{d}-2\exp(-C_{2}n)$,
$\dfrac{\|\hat{A}_{S}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\leq C_{3}\left\{\dfrac{\mu^{2}\max(a^{2},\sigma^{2})rM\log d}{n},a^{2}\mu\sqrt{\frac{\log d}{n}}\right\},$
provided that $n\geq 8C_{2}\mu rM\log d$. Here $C_{1},C_{3}$ depend on $L_{1},L_{2}$ and $C_{2}$ depends on $L_{1}$, where $L_{1},L_{2}$ are constants from Assumptions 1-2.
Comparing the convergence rate in Theorem 4 and the lower bound in (4), they are not matched due to the existence of the $\log d$ factor. Our new bound in (9) resolves the issue and shows that the estimator (8) is indeed minimax rate optimal. Regarding more comparisons between Theorems 3 and 4, we refer to remarks after Theorem 2 for similar comments.

### v3

@@ Abstract
In this paper, we demonstrate how a class of advanced matrix concentration inequalities, introduced in Brailovskaya and van Handel [2024], can be used to eliminate the dimensional factor in the convergence rate of matrix completion. This dimensional factor represents a significant gap between the upper bound and the minimax lower bound, especially in high dimension. Through a more precise spectral norm analysis, we remove the dimensional factors for three popular matrix completion estimators, thereby establishing their minimax rate optimality.
@@ 1 Introduction
In particular, many studies focus on the convergence rate of the nuclear norm penalization methods Koltchinskii et al. [2011], Negahban and Wainwright [2012], Klopp [2014], Yu et al. [2024]. However, the minimax lower bound, obtained in Koltchinskii et al. [2011] and Negahban and Wainwright [2012], does not include a logarithmic dimension factor that always appears in the upper bound of the convergence rate. This discrepancy between the two bounds becomes more pronounced in high-dimensional settings, highlighting a potential gap in the existing theoretical understanding. Due to the ubiquitousness of this issue, all the previous works have had to qualify their results by stating “our estimator is minimax optimal, up to a logarithmic factor”. In this paper, we leverage a class of sharp matrix concentration inequalities to remove the logarithmic dimension factor. After our work, the qualification will no longer be necessary.
@@
Let $A_{0}\in\mathbb{R}^{m_{1}\times m_{2}}$ be the unknown matrix we aim to recover. We observe independent samples $\{(X_{i},Y_{i})\}_{i=1}^{n}$ from the model
$Y_{i}=\langle X_{i},A_{0}\rangle+\xi_{i},\quad i=1,\ldots,n,$ (1)
where $X_{i}\in\mathbb{R}^{m_{1}\times m_{2}}$, $\langle X_{i},A_{0}\rangle=\text{tr}(X_{i}^{T}A_{0})$, and $\text{tr}(B)$ denotes the trace of matrix $B$. The sampling matrices $X_{i}\in\mathbb{R}^{m_{1}\times m_{2}}$ take values from $\{e_{j}(m_{1})e_{k}^{T}(m_{2}),1\leq j\leq m_{1},1\leq k\leq m_{2}\}$, where $\{e_{j}(m_{1})\}_{j=1}^{m_{1}},\{e_{k}(m_{2})\}_{k=1}^{m_{2}}$ denote standard basis vectors in $\mathbb{R}^{m_{1}},\mathbb{R}^{m_{2}}$, respectively. This implies that each $X_{i}$ only has one non-zero entry, indicating the location of the sampled entry.
@@
$P_{jk}=\mathbb{P}(X_{i}=e_{j}(m_{1})e_{k}^{T}(m_{2})),\quad 1\leq j\leq m_{1},1\leq k\leq m_{2}.$
$X_{i}$ and $\xi_{i}$ can be dependent. Throughout the paper, we consider that $\text{rank}(A_{0})\leq r$ and $\|A_{0}\|_{\infty}\leq a$, where $\|A\|_{\infty}=\max_{1\leq j\leq m_{1},1\leq k\leq m_{2}}|a_{jk}|$ is the entrywise maximum norm of $A=(a_{jk})$.
@@
The matrix completion problem is often addressed by a penalized regression. Assuming $A_{0}$ is of low rank, a natural choice for the penalty is the rank of a matrix. However, the rank penalty is non-convex and makes the corresponding program difficult to solve. The nuclear norm $\|\cdot\|_{*}$, namely the sum of all the singular values of a matrix, is usually used as a convex relaxation for the rank. There are many classical works on nuclear norm penalization methods under different assumptions. Assuming the noise has exponential decay, Koltchinskii et al. [2011] consider the case that the sampling distribution is known; Negahban and Wainwright [2012] study the situation where the row indices and the column indices of the observations are sampled independently and the penalty is a weighte nuclear norm; Klopp [2014] explores the problem for general sampling distributions. Consider the classical nuclear norm penalized least squares estimator
$\hat{A}=\operatorname*{argmin}_{\|A\|_{\infty}\leq a}\left\{\dfrac{1}{n}\sum_{i=1}^{n}(Y_{i}-\langle X_{i},A\rangle)^{2}+\lambda\|A\|_{*}\right\}.$
@@
$\dfrac{\|\hat{A}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\apprle\log(m_{1}+m_{2})\dfrac{r\max(m_{1},m_{2})}{n},$ (2)
@@
$\inf_{\hat{A}}\sup_{\mathcal{A}}\mathbb{P}\left(\dfrac{\|\hat{A}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\apprge\dfrac{r\max(m_{1},m_{2})}{n}\right)\geq C,$
@@
By our analysis, this gap arises essentially from the spectral norm analysis of some random matrices. In matrix completion, we frequently need to control the spectral norms of random matrices of the form $\frac{1}{n}\sum_{i=1}^{n}\zeta_{i}X_{i}$, where $\zeta_{i}$ are bounded or light-tailed random variables. The use of traditional concentration inequalities in this context typically introduces a logarithmic dependence on the dimension. This limitation is not unique to matrix completion—it arises in a variety of statistical and mathematical problems and has garnered significant attention in recent years. Notably, Bandeira et al. [2023] and Brailovskaya and van Handel [2024] proposed sharp new concentration inequalities that can successfully remove the logarithmic dependence in various settings. These results will serve as key tools in our paper.
Another point worth mentioning is that there is an alternative sampling model in matrix completion. In this model, there are Bernoulli random variables $\eta_{ij}$ indicating whether the entry of $A_{0}$ at location $(i,j)$ is observed. Under this assumption, the observations can be written as:
[…]
where $\xi_{ij}$ represents the noise at location $(i,j)$ and $A_{0}=(a_{0,ij})$. This sampling model is often referred to as sampling without replacement because each entry can be selected at most once. A rich body of literature has investigated this model [Candes and Plan, 2010, Chen and Wainwright, 2015, Candes and Recht, 2012, Keshavan et al., 2009, Chen and Wainwright, 2015, Chen et al., 2020]. As a contrast, the sampling strategy considered in our paper is known as sampling with replacement, as it allows repeated observations of the same entry. Since $n$ is often assumed to be much smaller than the matrix dimension $m_{1}\times m_{2}$, the number of repeated observations could be very small. Despite different sampling strategies, the theoretical results derived in both models are often compared side by side [Chen and Wainwright, 2015, Chen et al., 2020]. For instance, some works derive bounds of the following form [Chen and Wainwright, 2015]:
$\quad\frac{\|\widehat{A}-A_{0}\|_{F}^{2}}{m_{1}m_{2}}\lesssim\frac{r\max(m_{1},m_{2})}{m_{1}m_{2}p},$ (3)
under the non-replacement model with $\eta_{ij}\sim\text{Bernoulli}(p)$. Here, as the expected sample size, $m_{1}m_{2}p$ plays a role as $n$ in the with-replacement model. Compared to (3), the extra $\log(m_{1}+m_{2})$ factor in (2) is a common critique of algorithms developed under the with-replacement model[Chen and Wainwright, 2015, Chen et al., 2020]. Therefore, removing this logarithmic factor can help significantly defend the theoretical validity of the algorithms [Koltchinskii et al., 2011, Negahban and Wainwright, 2012, Klopp, 2014].
@@
In this paper, we revisit three estimators that have been studied in other research works:
  • The unknown matrix is of low rank and the noise is heavy tailed. The estimator in Yu et al. [2024] is designed for the case that the noise is only assumed to have finite second moment.
  • The unknown matrix is of low rank and the noise is sub-Gaussian. Nuclear norm penalized least squares is a common approach under this setting [Koltchinskii et al., 2011, Negahban and Wainwright, 2012, Klopp, 2014]. We consider the estimator proposed in Klopp [2014] to illustrate our new results.
  • The unknown matrix is of low rank and the noise is sub-Gaussian with unknown variance. The parameter tuning in matrix completion often requires prior knowledge of the noise variance. Klopp [2014] proposes an estimator to solve such a problem.
All the previous results on the convergence rates of the above three estimators involve the logarithmic dimension factor. We remove the dimensional factor for all the three estimators, thereby proving their minimax rate optimality.
The remainder of the paper is organized as follows. Continuing in this section, we introduce the notation. In Section 2, we present our main results. We give some concluding remarks in Section 3. Section 4 is dedicated to the proofs.
@@
To present our results, we collect the notations used throughout this paper for reader’s convenience.
@@
  • Let $A=(a_{ij})$, for $1\leq i\leq m_{1},1\leq j\leq m_{2}$. Let $\{\sigma_{l}\}_{l=1}^{m}$ denote ordered singular values of $A$: $\sigma_{1}\geq\sigma_{2}\geq\cdots\geq\sigma_{m}$. $\|A\|_{F}=\sqrt{\sum_{i=1}^{m_{1}}\sum_{j=1}^{m_{2}}a_{ij}^{2}}$, $\|A\|_{*}=\sum_{l=1}^{m}\sigma_{l}$, $\|A\|_{\infty}=\max_{i\leq m_{1},j\leq m_{2}}|a_{ij}|$, $\|A\|=\sigma_{1}$. $\|A\|_{w(F)}=\sqrt{\sum_{i=1}^{m_{1}}\sum_{i=1}^{m_{2}}a^{2}_{ij}P_{ij}}$ is a weighted Frobenius norm associated with the distribution of the sampling matrices $X_{i}$.
@@
    $\mathbb{P}(\epsilon_{i}=1)=\mathbb{P}(\epsilon_{i}=-1)=\frac{1}{2}.$
    They are independent of the observations $\{(X_{i},Y_{i})\}_{i=1}^{n}$.
  • We write $(\log d)^{\alpha}$ as $\log^{\alpha}d$ and $(\mathbb{E}[X])^{\alpha}$ as $\mathbb{E}^{\alpha}[X]$ for aesthetic reasons.
  • A random variable $X$ is called $\psi_{\alpha}$ random variable if it has finite $\psi_{\alpha}$ norm,
    $\|X\|_{\psi_{\alpha}}=\inf\left\{C>0;\mathbb{E}\exp\left(\left(\frac{|X|}{C}\right)^{\alpha}\right)\leq 2\right\}.$
    In particular, when $\alpha=2$, $\|\cdot\|_{\psi_{2}}$ is called sub-gaussian norm and X is called sub-gaussian random variable.
@@ 2 Main Results
In this section, we present new upper bounds for the three estimators mentioned in Section 1. In particular, compared to the existing bounds, the new bounds do not involve the dimension factor $\log d$ and hence match with the minimax lower bounds.
At first, we introduce assumptions used in our paper. Regarding the noise $\xi_{i}$, they can be heavy tailed or sub-Gaussian in different contexts and their distributions can depend on the sampled entry. We introduce the following assumptions.
@@
Given $X_{i}$, the noise variables satisfy
$E[\xi_{i}|X_{i}]=0$
and their second moments are uniformly bounded
$E[\xi_{i}^{2}|X_{i}]\leq\sigma^{2}.$
Assumption 2.
The noise variables $\xi_{i}$ satisfy $E[\xi_{i}|X_{i}]=0$, $E[\xi^{2}_{i}]=\sigma^{2}$ and given $X_{i}$, it holds uniformly
$E\exp(t\xi_{i}/\sigma)\leq\exp(L_{1}^{2}t^{2}),\forall t\in\mathbb{R},$
where $L_{1}>0$ is a constant.
@@
In following subsections, we introduce our $\log d$-free error bounds.